4 cubic metres of tin are extended by hammering the metal so as to cover an area of 1.6 deca ares. Find the thickness of the tin sheet?
Answer by mananth(16946)

1 deca ares = 1000m^2
1.6 deca ares = 1600 m^2
volume of sheet = area *thickness
4m^3 =1600 *thickness
4/1600 = thickness
=1/400
=0.0025 m
=2.5 mm